# Emergent synchrony in oscillator networks with adaptive arbitrary-order interactions

### v1

@@ Abstract
Acknowledgements — The authors acknowledge the financial support provided via NBRC Core Funds. DB acknowledges discussions with Dr. Proloy Das (NBRC), PV Vinitha (NBRC), and Shashwati Tripathi (NBRC).
@@
Declaration of competing interest — The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Appendix
Varying system size — In Fig. 5, the figures highlight the characteristic change in the order parameter dynamics as the system size $N$ is increased for different values of the pairwise coupling strength $\epsilon_{1}$.
The results demonstrate that, for values of $\epsilon_{1}$ given by $0.6$, $1.3$, and $1.8$, the blue curve (asynchronous initial conditions) loses its bi-modal structure and becomes sharply peaked near $r=0$ as the value of $N$ is increased; see Fig. 5(a)-5(i). In contrast, for $\epsilon_{1}=2.6$, the blue curve becomes sharply peaked near $r\approx 0.8$ upon increasing $N$; see Figs. 5(j)-5(l). These observations are consistent with the fact that, in the limit $N\rightarrow\infty$, the system has a stable fixed point at $r=0$ $\forall$ $\epsilon_{1}\leq 2$. The observations regarding the red curve (synchronous initial conditions) are markedly different. In particular, for $\epsilon_{1}=0.6$, the red curve completely loses its bi-modal structure and becomes nearly identical to the blue curve with increasing $N$. This is consistent with previous analysis, which showed that $r=0$ is the only steady synchronous state for this parameter value in the case $N\rightarrow\infty$. However, upon increasing $N$ for $\epsilon_{1}=1.3$, the red curve becomes locally bi-modal around $r\approx 0.8$. This indicates the presence of oscillatory dynamics in the order parameter and is consistent with the previous findings. Finally, upon increasing $\epsilon_{1}$ to $1.8$ and $2.6$, the red curve becomes unimodal with a sharp peak near $r\approx 0.8$ as the value of $N$ is increased, which is indicative of a stable synchronous fixed point.
@@
Simulation details — A direct simulation of the Kuramoto model with a coupling term of the form
@@
$X_{d}=\sum_{\mathclap{j_{1},..,j_{d}=1}}^{N}\sin\left(\sum_{k=1}^{d}c_{k}\phi_{j_{k}}\right),\ Y_{d}=\sum_{\mathclap{j_{1},..,j_{d}=1}}^{N}\cos\left(\sum_{k=1}^{d}c_{k}\phi_{j_{k}}\right).$ (12)
@@
$C_{d}=\sum_{j=1}^{N}\cos(c_{d}\phi_{j}),\ S_{d}=\sum_{j=1}^{N}\sin(c_{d}\phi_{j}).$ (14)
Similarly, $Y_{d}=Y_{d-1}C_{d}-X_{d-1}S_{d}$, whereas $(X_{1},Y_{1})=(S_{1},C_{1})$. This provides an iterative method to efficiently compute the interaction terms of the form of Eq. (11).

### v2

@@ Acknowledgements
The authors acknowledge the financial support provided via NBRC Core Funds. DB acknowledges discussions with Dr. Proloy Das (NBRC), PV Vinitha (NBRC), and Shashwati Tripathi (NBRC).
Declaration of competing interest
@@
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data Availability Statement
The data used in this article is available online [karmelic2022emergent, bougou2025mesoscale, dataset].
Appendix A Expression of Jacobians
The stability of the steady states, denoted by $(r_{s},\psi_{s})$, of Eqs. (14) and (15) can be determined from the eigenvalues of their Jacobian $J$, the elements $J_{ij}$ of which are given by
$\begin{split}J_{11}={}&\frac{-2r_{s}^{2}\Delta}{(1-r_{s}^{2})}+r_{s}(1-r_{s}^{2})\left[\sum_{d=1}^{D}\left.\frac{{\rm d}\sigma_{d}(r)}{{\rm d}r}\right|_{r_{s}}\cos(\delta_{d})\right.+\\ &\left.\sum_{d=2}^{D}\left.\frac{{\rm d}\sigma_{d}^{\prime}(r)}{{\rm d}r}\right|_{r_{s}}\cos(\chi_{d,s}^{\prime})\right],\end{split}$ (22)
$\begin{split}J_{21}={}&\frac{-2r_{s}\omega_{0}}{(1+r_{s}^{2})}-(1+r_{s}^{2})\left[\sum_{d=1}^{D}\left.\frac{{\rm d}\sigma_{d}(r)}{{\rm d}r}\right|_{r_{s}}\sin(\delta_{d})\right.+\\ &\left.\sum_{d=2}^{D}\left.\frac{{\rm d}\sigma_{d}^{\prime}(r)}{{\rm d}r}\right|_{r_{s}}\sin(\chi_{d,s}^{\prime})\right],\end{split}$ (23)
$J_{12}={}$ $2(h-1)r_{s}(1-r_{s}^{2})\sum_{d=2}^{D}\sigma_{d}^{\prime}(r_{s})\sin(\chi^{\prime}_{d,s}),$ (24)
and
$J_{22}={}$ $2(h-1)(1+r_{s}^{2})\sum_{d=2}^{D}\sigma_{d}^{\prime}(r_{s})\cos(\chi_{d,s}^{\prime}).$ (25)
For the choice of parameter values of the simplified model in Sec. 3, the expressions for the elements of the Jacobian reduce to the form
$\begin{split}J_{11}={}&\frac{-2r_{s}^{2}\Delta}{(1-r_{s}^{2})}+r_{s}(1-r_{s}^{2})\left[\left.\frac{{\rm d}\sigma_{1}(r)}{{\rm d}r}\right|_{r_{s}}+\right.\\ &\left.\frac{1}{\sigma_{2}(r_{s})}\left(\frac{\Delta}{1-r_{s}^{2}}-\sigma_{1}(r_{s})\right)\left.\frac{{\rm d}\sigma_{2}(r)}{{\rm d}r}\right|_{r_{s}}\right],\end{split}$ (26)
$J_{21}={}$ $\frac{-\omega_{0}}{(1+r_{s}^{2})}\left[2r_{s}+\frac{1+r_{s}^{2}}{\sigma_{2}(r_{s})}\left.\frac{{\rm d}\sigma_{2}(r)}{{\rm d}r}\right|_{r_{s}}\right],$ (27)
$J_{12}={}$ $2r_{s}(1-r_{s}^{2})\left[\frac{\omega_{0}}{1+r_{s}^{2}}\right],$ (28)
and
$J_{22}={}$ $2(1+r_{s}^{2})\left[\frac{\Delta}{1-r_{s}^{2}}-\Big(\sigma_{1}(r_{s})+\sigma_{2}(r_{s})\Big)\right],$ (29)
where
$\sigma_{1}(r)=\frac{\epsilon_{1}}{2\left(1+\gamma r\right)^{2}},\quad\sigma_{2}(r)=\frac{\epsilon_{2}}{2}(1+\gamma r)^{2}r^{2},$ (30)
and
$\frac{{\rm d}\sigma_{1}(r)}{{\rm d}r}=\frac{-\gamma\epsilon_{1}}{(1+\gamma r)^{3}},\quad\frac{{\rm d}\sigma_{2}(r)}{{\rm d}r}=\epsilon_{2}r\left[1+\gamma r(3+2\gamma r)\right],$ (31)
respectively. Note that the expressions depend solely on $r_{s}$ in this case, thus eliminating the need to separately calculate $\psi_{s}$, thereby simplifying calculations.
Appendix B Simulation details
@@
A direct simulation of the Kuramoto model, such as those with repeated initial conditions and different parameter values (i.e., Figs. 4 and 5), with a coupling term of the form
@@
$X_{d}=\sum_{\mathclap{j_{1},..,j_{d}=1}}^{N}\sin\left(\sum_{k=1}^{d}c_{k}\phi_{j_{k}}\right),\quad Y_{d}=\sum_{\mathclap{j_{1},..,j_{d}=1}}^{N}\cos\left(\sum_{k=1}^{d}c_{k}\phi_{j_{k}}\right).$ (33)
@@
$C_{d}=\sum_{j=1}^{N}\cos(c_{d}\phi_{j}),\quad S_{d}=\sum_{j=1}^{N}\sin(c_{d}\phi_{j}).$ (35)
Similarly, $Y_{d}=Y_{d-1}C_{d}-X_{d-1}S_{d}$, whereas $(X_{1},Y_{1})=(S_{1},C_{1})$. This provides an iterative method to efficiently compute the interaction terms of the form of Eq. (32), and can be extended to Kuramoto-like coupling terms in any dynamical systems.
Unless mentioned otherwise, the simulations of Eq. (18) were performed using the popular 4th-order Runge-Kutta algorithm [butcher1996history], for a system of $N=10^{4}$ oscillators for $10^{4}$ time-steps, with each individual time-step being $10^{-2}$ seconds. The algorithm was implemented in C++, using OpenMP for parallelisation of the computationally intensive portions [dagum1998openmp]. The solutions of Eq. (21) were estimated using the function fsolve(), whereas eigenvalues of the Jacobians in A were obtained using the function eig() in MATLAB. All the computations were executed on a workstation equipped with an Intel Core i9-14900K processor and 64 GB DDR5 RAM.
Appendix C Varying system size
In Fig. 6, the figures highlight the characteristic change in the order parameter dynamics as the system size $N$ is increased for different values of the pairwise coupling strength $\epsilon_{1}$.
The results demonstrate that, for values of $\epsilon_{1}$ given by $0.6$, $1.3$, and $1.8$, the blue curve (asynchronous initial conditions) loses its bi-modal structure and becomes sharply peaked near $r=0$ as the value of $N$ is increased; see Fig. 6(a)-6(i). In contrast, for $\epsilon_{1}=2.6$, the blue curve becomes sharply peaked near $r\approx 0.8$ upon increasing $N$; see Figs. 6(j)-6(l). These observations are consistent with the fact that, in the limit $N\rightarrow\infty$, the system has a stable fixed point at $r=0$ $\forall$ $\epsilon_{1}\leq 2$. The observations regarding the red curve (synchronous initial conditions) are markedly different. In particular, for $\epsilon_{1}=0.6$, the red curve completely loses its bi-modal structure and becomes nearly identical to the blue curve with increasing $N$. This is consistent with previous analysis, which showed that $r=0$ is the only steady synchronous state for this parameter value in the case $N\rightarrow\infty$. However, upon increasing $N$ for $\epsilon_{1}=1.3$, the red curve becomes locally bi-modal around $r\approx 0.8$. This indicates the presence of oscillatory dynamics in the order parameter and is consistent with the previous findings. Finally, upon increasing $\epsilon_{1}$ to $1.8$ and $2.6$, the red curve becomes unimodal with a sharp peak near $r\approx 0.8$ as the value of $N$ is increased, which is indicative of a stable synchronous fixed point.